1. Which one of these in not a linear equation?
0.07x - 0.09x = 0.57
5t - 11t = -6t
7x + 9(x-2)= -5x
6y(2 squared)-3y + 1 = 0
2. Find the indicated intersection or union.
{q, s, u, v, w, x, y, z} u (upside down) {s}
3. 1 + 1
- - = c, for b (solve for b)
a b
4. Mark swims at a speed of 5.8 mph in still water. The lake he's in flows at a speed of 3.7 mph. How long will it take Mark to swim 2.2 mi downstream? Round to the nearest tenth of an hour, if applicable.

Answer by user_dude2008(1862)

1. 6y(2 squared)-3y + 1 = 0 is not linear
2. {q, s, u, v, w, x, y, z} u (upside down) {s} = {s}
3. 1/a+1/b=c ----> Solution: b = a/(ac-1)